OpenStudy (anonymous):

A 30°-60°-90° triangle has a hypotenuse with a length of 10. What is the length of the longer leg of the triangle?

OpenStudy (anonymous):

10/2 = 5 shortest leg of a 306090 triangle is always half the hypotenuse also cos(30)=x/10 so 10cos(30) = x cos (30) = 1/2 so 10 *1/2 = 5 sin(30) = square root of 3 (read: root 3 or 3 ^.5) over two so if you want the other leg (the longer short one) then you do 10(3^.5)/2 or 5(3^.5) are you in CAVA?

OpenStudy (anonymous):

Thank you! And no, i'm not in CAVA.

OpenStudy (anonymous):

You're Welcome :) Oh, I had the same question a while ago and thanks for the medal :D